OpenStudy (jake124):

which number is the solution of m + 91 = 157 A. 58 B. 66 my answer C. 68 D. 248 which number is the solution of n/3 = -12 A. -36 my answer B. -4 C. 4 D. 36 use mental math to find the solution of 19b = 190 A. 10 my answer B. 11 C. 171 D. 208 please check my answers

OpenStudy (jake124):

estimate the solution of y + 1.8 = 3.14 to the nearest whole number A.0 B. 1 C. 2 my answer D. 5

DivineSolar (divinesolar):

Alright I will check them.

DivineSolar (divinesolar):

m + 91 - 91 = 157 - 91 > 66 > Correct.

OpenStudy (retireed):

Is 2.8 closer to 3.14 or is 3.8 closer to 3.14?

DivineSolar (divinesolar):

n/3 * 3 = -12 * 3 > -36 > Correct

DivineSolar (divinesolar):

19b/19 = 190/19 > 10 > Correct

DivineSolar (divinesolar):

The 4th one also looks correct.

OpenStudy (retireed):

Is 2.8 closer to 3.14 or is 3.8 closer to 3.14?

DivineSolar (divinesolar):

3.8 > 4 2.8 > 3 < 2.8 is closer to 3.14

OpenStudy (jake124):

so 2 is correct

DivineSolar (divinesolar):

Correct.

OpenStudy (retireed):

No One is correct the answer is B for the fourth question.

OpenStudy (jake124):

thank u it was 1

DivineSolar (divinesolar):

Oh yes my bad you get 1.3 when doing this equation. Oopsy! So yes A is correct.

DivineSolar (divinesolar):

Or not A but B

DivineSolar (divinesolar):

Thought A was 1 lol.

DivineSolar (divinesolar):

1.3 > 1

OpenStudy (phi):

estimate the solution of y + 1.8 = 3.14 to the nearest whole number they want you to change that to y+2 = 3 and then find y=1

OpenStudy (retireed):

1 + 1.8 = 2.8 approximately 3 2 + 1.8 = 3.8 approximately 4 3.14 is approximately 3 One is the right Answer B.